OpenStudy (anonymous):

The following table shows the data of a random sample of fish collected from and later released into a lake: How many bass are estimated to be present if a random sample of 1,500 fish is collected from the lake? Type of Fish Bass Catfish Warmouth Number of fish 60 70 20

OpenStudy (danjs):

(sample size) * (# bass) / (Total Fish)

OpenStudy (danjs):

Take the fraction that are bass, and multiply that by your sample size

OpenStudy (anonymous):

60/1500?

OpenStudy (danjs):

Ok, in the sample, there are 60 bass and (60+70+20) = 150 total fish

OpenStudy (danjs):

the fraction that are bass is 60/150

OpenStudy (danjs):

or 2/5 or 40%

OpenStudy (anonymous):

then add the total after?

OpenStudy (danjs):

If 1500 Fish are sampled, you would expect 40% to be bass

OpenStudy (danjs):

1500*0.4 = 600 estimated bass

OpenStudy (danjs):

This is what we did \[\frac{ 60 }{ 60+70+20 } = \frac{ x }{ 1500 }\]

OpenStudy (danjs):

Just scaled up the sample size to 1500 from 150 with the same expectations on percent of bass

OpenStudy (anonymous):

@DanJS how would I come to the total

OpenStudy (anonymous):

600 would be the answer